if three cans of corn cost you $8.00 how much will one gross cost you?
what dose gross mean?

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(2446)

gross = twelve dozen = 12*12 = 144
.
if three cans of corn cost you $8.00 how much will one gross cost you?
8/3 * 144
= 8*48
= $384

Answer by Alan3354(6092)

if three cans of corn cost you $8.00 how much will one gross cost you?
what dose gross mean?
-------------
A gross is 12 dozen, = 144, continuing the American tradition of using tedious units of measure, like inches, feet, yards, ounces, fluid ounces, pounds, etc.
------------
Anyway, at $8 for 3, 144 is 48 times as many, so they'll cost
$8 times 48
= $384
------------
You should ask for a quantity discount.